The function s(m) = 5 + 4(m – 1) represents the number of stamps in Rudy’s stamp collection in month m.
What does the value 4 represent in this situation?
(Points : 1)
a. Rudy’s collection is worth $4.
b. Rudy collected stamps for 4 months.
c. Rudy began his collection with 4 stamps.
d. Rudy adds 4 more stamps to his collection every month.

Answer by ewatrrr(24785)

s(m) = 5 + 4(m – 1)
....
1st month
s(1) = 5 + 4(0)
s(2) = 5 +4(1)
s(3) = 5 + 4(2)
etc
....
d. Rudy adds 4 more stamps to his collection every month.